# Exactly solvable dissipative dynamics and one-form strong-to-weak spontaneous symmetry breaking in interacting two-dimensional spin systems

### v1

@@ IV.5 Steady states and excitations from strings
    Open strings $W_{\mathcal{P}}$ ($\mu,\nu=0$). This final type of string carries instead itinerant Majorana fermions $c_{j}$ at its endpoints, see Fig. 4(c). By the arguments laid out after Eq. (9), $W_{\mathcal{P}}$ does not commute with every $K_{ij}$, but it does commute with all $W_{\mathcal{C}}$ and $W_{\mathcal{P}}^{\mathrm{ch}}$. Hence, $s_{p}=s_{\ell}=\mathfrak{s}_{q}^{L}=\mathfrak{s}_{q}^{R}=1$ and there are no gauge excitations. $\sigma=W_{\mathcal{P}}\rho_{0}=\rho_{0}W_{\mathcal{P}}$ [^9] At the vectorized level, the equality of left and right actions translates to Eq. (44). is a decaying operator in the same symmetry sector as $\rho_{0}$ and is vectorized to $c_{i}c_{j}|0\rangle_{\bm{w}=\bm{1},\bm{\mathfrak{w}}=\bm{1}}=f^{\dagger}_{i}f^{\dagger}_{j}|0\rangle_{\bm{w}=\bm{1},\bm{\mathfrak{w}}=\bm{1}}$ [where we used Eqs. (33) and (44)], creating a superposition of fermionic excitations. On the other hand, the adjoint action of $W_{p}$ leaves $\rho_{0}$ invariant, which in the vectorized picture can be understood as inserting both a $c_{j}$ and a $\tilde{c}_{j}$ fermion into the system, using the vacuum to reflect the $\tilde{c}_{j}$ to another $c_{j}$ [Eq. (27)], and then annihilating it with the first $c_{j}$.
More general strings, not of the form (6) can also be considered but lead to no new excitations. An example, closed strings that are not the products of $K_{ij}$ but some other gamma-matrix bilinears, do not commute with the fluxes and thus induce vortices along the body of the string. Composing the previous operations, every operator in any symmetry sector can be reached, thus completing the enumeration of possible excitations in both the gamma-matrix and fermion representations.
@@ V.1 Diagonalization of the Lindbladian
[…]
@@
referred to as normal master modes (NMM) [76], satisfy
@@ V.2 Parity of physical states from the Pfaffian
A state with a total of $\nu=\sum_{k}\nu_{k}$ populated NMM is physical if and only if
@@
Moreover, if there are $|\mathfrak{I}|$ purely imaginary $\beta_{k}$, then $\det O=\mathrm{sgn}\,\big{(}(-\mathrm{i})^{|\mathfrak{I}|}\mathrm{Pf}\,\mathrm{i}A\big{)}$, where $\mathrm{Pf}\,$ is the Pfaffian.
@@
Performing the change of basis between the original and rotated Majoranas, we have that $(-1)^{\mathcal{N}_{f}}=\prod_{j=1}^{N}(-\mathrm{i}c_{j}\tilde{c}_{j})=(-\mathrm{i})^{N}\sum_{j_{1}\cdots j_{2N}}O_{j_{1}1}\cdots O_{j_{2N}2N}a_{j_{1}}\cdots a_{j_{2N}}$. Terms with any repeated indices do not contribute to the sum: if $j_{a}=j_{b}$ for some $a\neq b$, then $\sum_{j_{a}}O_{j_{a}a}O_{j_{a}b}a_{j_{a}}^{2}=(OO^{T})_{ab}=\delta_{ab}=0$. Thus, only the terms with all indices $j_{1},\dots,j_{N}$ distinct contribute. Since the product $a_{j_{1}}\cdots a_{j_{2N}}$ is fully antisymmetric in its indices, we can replace it with $\varepsilon_{j_{1}\cdots j_{2N}}a_{1}\cdots a_{2N}$, where $\varepsilon_{j_{1}\cdots j_{2N}}$ is the fully antisymmetric symbol, and it follows that $(-1)^{\mathcal{N}_{f}}=(-1)^{\mathcal{N}_{d}}\sum_{j_{1}\cdots j_{2N}}\varepsilon_{j_{1}\cdots j_{2N}}O_{j_{1}1}\cdots O_{j_{2N}2N}=(-1)^{\mathcal{N}_{d}}\det O$, where we used Eq. (59), i.e., the change of basis is parity preserving if and only if $O$ is smoothly connected to the identity, which is parity preserving. Using Eq. (41), Eq. (60) follows. Furthermore, $\mathrm{Pf}\,\mathrm{i}A=\mathrm{Pf}\,(O^{T}\mathrm{i}BO)=\det O\,\mathrm{Pf}\,\mathrm{i}B=\det O\prod_{j=1}^{N}\beta_{j}$. Since $A$ is similar to a real matrix ($A=TA^{*}T^{-1}$ with $T=P(\sigma^{z}\otimes\mathbbm{1})P^{T}$), and ±$\pm\beta_{j}$ are its eigenvalues, we have three possibilities: (i) $\beta_{k}\in\mathbb{R}$ and it is, by assumption, positive; (ii) $\beta_{k}$ come in complex-conjugate pairs, i.e., there exists some $k\neq j$ such that $\beta_{k}=\beta_{j}^{*}\implies\beta_{j}\beta_{k}>0$; or (iii) $\beta_{k}\in\mathrm{\mathrm{i}}\mathbb{R}$ with, by assumption, $\mathrm{Im}\beta_{k}>0$. Hence, $\mathrm{Pf}\,\mathrm{i}B=\mathrm{i}^{|\mathfrak{I}|}R$ for some $R>0$. Since $O$ is orthogonal, ±$\pm 1=\det O=\mathrm{Pf}\,\mathrm{i}A/\mathrm{Pf}\,\mathrm{i}B=(-\mathrm{i})^{|\mathfrak{I}|}\mathrm{Pf}\,\mathrm{i}A/R$ and taking the sign of both sides, we arrive at the desired result, $\det O=\mathrm{sgn}\,\big{(}(-\mathrm{i})^{|\mathfrak{I}|}\mathrm{Pf}\,\mathrm{i}A\big{)}$. ∎
@@ V.3 Spectral gap and relaxation rates
Let us first consider steady-state sectors (i.e., $\mathbb{A}=\widetilde{\mathbb{A}}$, $\mathfrak{U}=\emptyset$, $\mathfrak{V}=\emptyset$). For $\gamma_{j}\to 0$, $\mathrm{Pf}\,\mathrm{i}A\to\mathrm{Pf}\,\mathrm{i}\mathbb{A}\mathrm{Pf}\,\mathrm{i}\widetilde{\mathbb{A}}=(\mathrm{Pf}\,\mathrm{i}\mathbb{A})^{2}=\det\mathrm{i}\mathbb{A}$ and $|\mathfrak{I}|\to N$ (because $A$ becomes real antisymmetric). Thus, $(-\mathrm{i})^{|\mathfrak{I}|}\mathrm{Pf}\,\mathrm{i}A\to\det\mathbb{A}\geq 0$ (since $\mathbb{A}$ is real antisymmetric), hence $\det O\to 1$. Then, by continuity of $\det$, the NMM vacuum $|\Omega\rangle$ is physical for small $\gamma_{j}$ and, by Eq. (58), it is the slowest decaying state. Hence, it must coincide with the steady state (i.e., the original adjoint-fermion vacuum $|0\rangle_{\bm{w},\bm{\mathfrak{w}}=\bm{1}}$) and we must have $B_{0}=0$, i.e., $\sum_{k}\beta_{k}/2=\sum_{j}\gamma_{j}$ [76] (this sum rule does not hold outside steady-state sectors). The existence of the vacuum steady state for any $\gamma_{j}$ indicates that the Pfaffian cannot change sign upon increasing $\gamma_{j}$ and all physical states in these sectors have an even number of fermion excitations, $\prod_{j}(-1)^{\nu_{j}}=1$. As discussed above, there is an additional U(1) symmetry in these sectors and excitations are labeled by the adjoint-fermion number $n$. The gap in this sector is given by $\Delta_{n}=\sum_{k=1}^{n}\mathrm{Re}\beta_{k}$ ($n\in\{0,2,4,\dots,N\}$). This case will be treated in detail in Sec. VI.
@@ V.4 Numerical results
In Fig. 5, we plot the gap $\Delta$ as a function of dissipation strength $\gamma$ for the honeycomb lattice on the torus for fixed $\mathfrak{V}=\emptyset$ and different choices of $\mathfrak{U}$. The qualitative features of the gap for the square [35], honeycomb (Fig. 5), and triangular (App. C) lattices are similar, indicating that dissipative dynamics are more strongly determined by the structure of the Lindbladian (50) than by the microscopic details of the couplings $J_{ij}$ (i.e., of the underlying graph).
@@
Some general features, which can be understood perturbatively in $\gamma$ or $1/\gamma$, are immediately visible at finite $N$: the gap grows linearly as $\gamma\to\infty$ and closes as $\gamma\to 0$. At intermediate values of $\gamma$, a rich nonmonotonic behavior occurs due to a succession of crossings of the single-particle eigenvalues $\mathrm{Re}\beta_{k}$ or exceptional points (i.e., values of $\gamma$ at which a complex-conjugate pair $(\beta_{k},\beta_{k}^{*})$ merges into a doubly-degenerate real eigenvalue). The kinks in the gap that result from these crossings and exceptional points are particularly visible, e.g., in Fig. 5(c).
@@
At small $\gamma$, the slope of the gap increases with increasing $N$, see Figs. 5(a)–(c), which suggests that the weak dissipation ($\gamma\to 0$) and thermodynamic ($N\to\infty$) limits do not commute. Hence, if the thermodynamic limit is taken first, the gap converges to a finite constant of order $J$ (where $J$ is the typical coupling strength of $H$) in the limit of $\gamma\to 0$, signaling a much faster relaxation than expected from perturbation theory. This phenomenon, dubbed anomalous relaxation [119], signals that internal interactions, not external dissipation, govern the relaxation. It has been recently observed in a variety of systems [118, 119, 120, 139, 140, 141, 35, 121, 122, 142, 111]. It is proposed to be a generic feature of ergodic many-body quantum dynamics [120, 121, 122, 143, 144] (via a relation to quantum Ruelle-Pollicott resonances [120, 122]). [^11] The noncommutativity of limits can be understood heuristically as follows [120]. For an $N$-site system with bulk dissipation $\gamma$, there is a natural effective dissipation strength $\gamma_{\mathrm{eff}}=N\gamma$. It is thus manifest that the two limits do not commute: $\lim_{N\to\infty}\lim_{\gamma\to 0}\gamma_{\mathrm{eff}}=0$, while $\lim_{\gamma\to 0}\lim_{N\to\infty}\gamma_{\mathrm{eff}}=\infty$. Nevertheless, the weak-dissipation dynamics of our gamma-matrix model are not fully mixing because of the emergent integrability arising from the fractionalizing into an exponentially large number of gauge sectors with free-fermion excitations. In every sector gauge-equivalent to $\mathfrak{U}=\emptyset$, the Linbdladian has an additional conserved quantity in the limit of $\gamma\to 0$, namely the open gauge strings $W^{\mathrm{ch}}_{\mathcal{P}}$ (together with $H$ itself, these exhaust the conservation laws of the Hamiltonian and hence the $\gamma\to 0$ limit of the Lindbladian). By virtue of these extra conservation laws, there is an infinite number (in the thermodynamic limit) of gaps that close and the system as a whole evades anomalous relaxation, see, e.g., Fig. 5(d). By choosing the gauge where the particle-number conservation is manifest, we can analytically bound the gap in different particle-number sectors and prove the absence of anomalous relaxation, see Sec. VI.2. Notwithstanding, the (also extensive) number of sectors with finite gaps further illustrates the ubiquity of anomalous relaxation as a generic phenomenon in strongly interacting many-body dynamics.
@@
On the other hand, certain symmetry sectors display the quantum Zeno effect—the closing of the spectral gap with increasing dissipation strength—which was reported for homogeneous dissipation on the square lattice in Ref. 35. For the gap to close in this limit, there must exist an operator $\Sigma_{R}$ that commutes with all jump operators $L_{j}\sim\Gamma_{j}^{\mathrm{ch}}$. Such operators exist in sectors that either are gauge-equivalent to $\mathfrak{U}=\emptyset$ or have $\mathfrak{V}=\emptyset$ and $(-1)^{|\mathfrak{U}|}=-1$. The former will be discussed in detail in Sec. VI.3 and here we focus on the latter. For any choice of $\mathfrak{U}$ with even $|\mathfrak{U}|$, we consider a product $\Sigma_{R}$ of open gauge strings $W_{\mathcal{P}}^{\mu\nu}$ ($\mu,\nu\in\{1,\dots,2k\}$) whose endpoints connect a pair of sites from which a gauge flip emanates, such that all flips are connected by one such string. From Sec. IV.5, $\sigma=\rho_{\bm{w}}\Sigma_{R}$ belongs to the sector labeled by $\mathfrak{U}$ and, because $W_{\mathcal{P}}^{\mu\nu}$ is built from two gamma matrices on each site, it commutes with all chiral elements $\Gamma_{j}^{\mathrm{ch}}$. Hence, $\sigma$ is annihilated by $\mathcal{L}$ in the limit $\gamma_{j}\to\infty$. On the other hand, if $|\mathfrak{U}|$ is odd, one of the gauge flips is not paired up. Hence, there exists a site $j$ for which $\Gamma_{j}^{\mu}$ appears only once in $\Sigma_{R}$, which thus anticommutes with $\Gamma_{j}^{\mathrm{ch}}$, and the gap diverges. These predictions are confirmed in Fig. 5.121212At first sight, this reasoning should imply that the gap diverges as $\gamma\to\infty$ in Fig. 5(d) because $|\mathfrak{U}|$ is odd. However, as discussed in Sec. IV.5, the action of $\Gamma_{j}^{\mathrm{ch}}$ is not to flip all the $\tilde{u}_{ij}$ around a site but to flip the corresponding $v_{j}$. A representative string in the gauge-equivalent representation with $\mathfrak{V}=\emptyset$ (which we are employing in the numerical calculations) is $\Sigma_{R}=D_{j}\Gamma_{j}^{\mathrm{ch}}$. This operator, besides flipping the intralayer gauge fields around site $j$, also anticommutes with the fermionic parity $(-1)^{\mathcal{N}_{f}}$, leading to the result of Fig. 5(d), which is also in agreement with following section, cf. Fig. 6.

### v2

@@ IV.5 Steady states and excitations from strings
    Open strings $W_{\mathcal{P}}$ ($\mu,\nu=0$). This final type of string carries instead itinerant Majorana fermions $c_{j}$ at its endpoints; see Fig. 4(c). By the arguments laid out after Eq. (9), $W_{\mathcal{P}}$ does not commute with every $K_{ij}$, but it does commute with all $W_{\mathcal{C}}$ and $W_{\mathcal{P}}^{\mathrm{ch}}$. Hence, $s_{p}=s_{\ell}=\mathfrak{s}_{q}^{L}=\mathfrak{s}_{q}^{R}=1$ and there are no gauge excitations. $\sigma=W_{\mathcal{P}}\rho_{0}=\rho_{0}W_{\mathcal{P}}$ [^9] At the vectorized level, the equality of left and right actions translates to Eq. (44). is a decaying operator in the same symmetry sector as $\rho_{0}$ and is vectorized to $c_{i}c_{j}|0\rangle_{\bm{w}=\bm{1},\bm{\mathfrak{w}}=\bm{1}}=f^{\dagger}_{i}f^{\dagger}_{j}|0\rangle_{\bm{w}=\bm{1},\bm{\mathfrak{w}}=\bm{1}}$ [where we used Eqs. (33) and (44)], creating a superposition of fermionic excitations. On the other hand, the adjoint action of $W_{p}$ leaves $\rho_{0}$ invariant, which in the vectorized picture can be understood as inserting both a $c_{j}$ and a $\tilde{c}_{j}$ fermion into the system, using the vacuum to reflect the $\tilde{c}_{j}$ to another $c_{j}$ [Eq. (27)], and then annihilating it with the first $c_{j}$.
More general strings, not of the form (6) can also be considered but lead to no new excitations. For example, closed strings that are not the products of $K_{ij}$ but some other gamma-matrix bilinears, do not commute with the fluxes and thus induce vortices along the body of the string. Composing the previous operations, every operator in any symmetry sector can be reached, thus completing the enumeration of possible excitations in both the gamma-matrix and fermion representations.
@@ V.1 Diagonalization of the Lindbladian
Because of the antisymmetry of $A$, its $2N$ eigenvalues come in pairs $(\beta_{k},-\beta_{k})$, where, without loss of generality, we order $0\leq\mathrm{Re}\beta_{1}\leq\cdots\leq\mathrm{Re}\beta_{N}$; if $\mathrm{Re}\beta_{k}=0$, then we take $\mathrm{Im}\beta_{k}>0$. [^10] Here, we assumed that the Lindbladian is sufficiently generic such that it is diagonalizable and that there are no eigenvalues $\beta_{k}=0$. The latter would imply there are multiple steady states in a single symmetry sector and lead to the breakdown of the Pfaffian method described below. Denoting the right eigenvectors of $A$ by $r$, $Ar_{2k-1}=\beta_{k}r_{2k-1}$ and $Ar_{2k}=-\beta_{k}r_{2k}$, we define the $2N\times 2N$ matrix $O$ whose rows are $O_{2k-1}=(r^{T}_{2k-1}+r^{T}_{2k})/\sqrt{2}$ and $O_{2k-1}=-\mathrm{i}(r^{T}_{2k-1}-r^{T}_{2k})/\sqrt{2}$. Since the eigenvectors of $A$ are complex and biorthogonal, $O$ is a complex orthogonal matrix, $OO^{T}=O^{T}O=\mathbbm{1}$, and we can use it to reduce $A$ to the canonical form $A=O^{T}BO$ [141, 77], with
@@
referred to as normal master modes (NMMs) [77], satisfy
@@ V.2 Parity of physical states from the Pfaffian
A state with a total of $\nu=\sum_{k}\nu_{k}$ populated NMMs is physical if and only if
@@
Moreover, if there are $|\mathfrak{I}|$ purely imaginary $\beta_{k}$, then
$\det O=\mathrm{sgn}\,\big((-\mathrm{i})^{|\mathfrak{I}|}\mathrm{Pf}\,\mathrm{i}A\big),$ (61)
where $\mathrm{Pf}\,$ is the Pfaffian.
@@
Performing the change of basis between the original and rotated Majoranas, we have that $(-1)^{\mathcal{N}_{f}}=\prod_{j=1}^{N}(-\mathrm{i}c_{j}\tilde{c}_{j})=(-\mathrm{i})^{N}\sum_{j_{1}\cdots j_{2N}}O_{j_{1}1}\cdots O_{j_{2N}2N}a_{j_{1}}\cdots a_{j_{2N}}$. Terms with any repeated indices do not contribute to the sum: if $j_{a}=j_{b}$ for some $a\neq b$, then $\sum_{j_{a}}O_{j_{a}a}O_{j_{a}b}a_{j_{a}}^{2}=(OO^{T})_{ab}=\delta_{ab}=0$. Thus, only the terms with all indices $j_{1},\dots,j_{N}$ distinct contribute. Since the product $a_{j_{1}}\cdots a_{j_{2N}}$ is fully antisymmetric in its indices, we can replace it with $\varepsilon_{j_{1}\cdots j_{2N}}a_{1}\cdots a_{2N}$, where $\varepsilon_{j_{1}\cdots j_{2N}}$ is the fully antisymmetric symbol, and it follows that $(-1)^{\mathcal{N}_{f}}=(-1)^{\mathcal{N}_{d}}\sum_{j_{1}\cdots j_{2N}}\varepsilon_{j_{1}\cdots j_{2N}}O_{j_{1}1}\cdots O_{j_{2N}2N}=(-1)^{\mathcal{N}_{d}}\det O$, where we used Eq. (59), i.e., the change of basis is parity preserving if and only if $O$ is smoothly connected to the identity, which is parity preserving. Using Eq. (41), Eq. (60) follows. Furthermore, $\mathrm{Pf}\,\mathrm{i}A=\mathrm{Pf}\,(O^{T}\mathrm{i}BO)=\det O\,\mathrm{Pf}\,\mathrm{i}B=\det O\prod_{j=1}^{N}\beta_{j}$. Since $A$ is similar to a real matrix ($A=TA^{*}T^{-1}$ with $T=P(\sigma^{z}\otimes\mathbbm{1})P^{T}$), and $\pm\beta_{j}$ are its eigenvalues, we have three possibilities: (i) $\beta_{k}\in\mathbb{R}$ and it is, by assumption, positive; (ii) $\beta_{k}$ come in complex-conjugate pairs, i.e., there exists some $k\neq j$ such that $\beta_{k}=\beta_{j}^{*}\implies\beta_{j}\beta_{k}>0$; or (iii) $\beta_{k}\in\mathrm{\mathrm{i}}\mathbb{R}$ with, by assumption, $\mathrm{Im}\beta_{k}>0$. Hence, $\mathrm{Pf}\,\mathrm{i}B=\mathrm{i}^{|\mathfrak{I}|}R$ for some $R>0$. Since $O$ is orthogonal, $\pm 1=\det O=\mathrm{Pf}\,\mathrm{i}A/\mathrm{Pf}\,\mathrm{i}B=(-\mathrm{i})^{|\mathfrak{I}|}\mathrm{Pf}\,\mathrm{i}A/R$, and taking the sign of both sides, we arrive at Eq. (61). ∎
@@ V.3 Spectral gap and relaxation rates
Figure 5: Spectral gap as a function of dissipation strength in fermion-parity-conserving sectors. The gap was computed for the honeycomb lattice with N=128N=128512512 and periodic boundary conditions on the torus (with the same number of sites along each direction) using the prescription of Sec. V.3, with the single-particle eigenvalues $\beta_{k}$ obtained from exact diagonalization of Eq. (52). We set uniform Hamiltonian couplings and gauge fields $J_{ij}=u_{ij}=1$ (which gives uniform 0-flux), uniform dissipation $\gamma_{i}=\gamma$ for all $i$, and fixed the gauge to $v_{i}=-1$ for all $i$ (i.e., $\mathfrak{V}=\emptyset$). Each panel shows a different type of intralayer gauge flips $\tilde{u}_{ij}=-1$ (i.e., choice of $\mathfrak{U}$): (a) a single arbitrary edge; (b) all edges of a given color; (c) all edges on a noncontractible loop; and (d) all edges emanating from an arbitrary vertex (which is gauge-equivalent to a single interlayer flip on that vertex and $\mathfrak{U}=\emptyset$, with conserved fermion number, cf. upper left panel of Fig. 6).
Let us first consider steady-state sectors (i.e., $\mathbb{A}=\widetilde{\mathbb{A}}$, $\mathfrak{U}=\emptyset$, $\mathfrak{V}=\emptyset$). For $\gamma_{j}\to 0$, $\mathrm{Pf}\,\mathrm{i}A\to\mathrm{Pf}\,\mathrm{i}\mathbb{A}\mathrm{Pf}\,\mathrm{i}\widetilde{\mathbb{A}}=(\mathrm{Pf}\,\mathrm{i}\mathbb{A})^{2}=\det\mathrm{i}\mathbb{A}$ and $|\mathfrak{I}|\to N$ (because $A$ becomes real antisymmetric). Thus, $(-\mathrm{i})^{|\mathfrak{I}|}\mathrm{Pf}\,\mathrm{i}A\to\det\mathbb{A}\geq 0$ (since $\mathbb{A}$ is real antisymmetric), hence $\det O\to 1$. Then, by continuity of $\det$, the NMM vacuum $|\Omega\rangle$ is physical for small $\gamma_{j}$ and, by Eq. (58), it is the slowest decaying state. Hence, it must coincide with the steady state (i.e., the original adjoint-fermion vacuum $|0\rangle_{\bm{w},\bm{\mathfrak{w}}=\bm{1}}$) and we must have $B_{0}=0$, i.e., $\sum_{k}\beta_{k}/2=\sum_{j}\gamma_{j}$ [77] (this sum rule does not hold outside steady-state sectors). The existence of the vacuum steady state for any $\gamma_{j}$ indicates that the Pfaffian cannot change sign upon increasing $\gamma_{j}$, and all physical states in these sectors have an even number of fermion excitations, $\prod_{j}(-1)^{\nu_{j}}=1$. As discussed above, there is an additional U(1) symmetry in these sectors, and excitations are labeled by the adjoint-fermion number $n$. The gap in this sector is given by $\Delta_{n}=\sum_{k=1}^{n}\mathrm{Re}\beta_{k}$ ($n\in\{0,2,4,\dots,N\}$). This case will be treated in detail in Sec. VI.
@@ V.4 Numerical results
In Fig. 5, we plot the gap $\Delta$ as a function of dissipation strength $\gamma$ for the honeycomb lattice on the torus for fixed $\mathfrak{V}=\emptyset$ and different choices of $\mathfrak{U}$. The qualitative features of the gap for the square [35], honeycomb (Fig. 5), and triangular (Appendix C) lattices are similar, indicating that dissipative dynamics are more strongly determined by the structure of the Lindbladian (50) than by the microscopic details of the couplings $J_{ij}$ (i.e., of the underlying graph).
@@
Some general features, which can be understood perturbatively in $\gamma$ or $1/\gamma$, are immediately visible at finite $N$: the gap grows linearly as $\gamma\to\infty$ and closes as $\gamma\to 0$. At intermediate values of $\gamma$, a rich nonmonotonic behavior occurs due to a succession of crossings of the single-particle eigenvalues $\mathrm{Re}\beta_{k}$ or exceptional points [i.e., values of $\gamma$ at which a complex-conjugate pair $(\beta_{k},\beta_{k}^{*})$ merges into a doubly-degenerate real eigenvalue]. The kinks in the gap that result from these crossings and exceptional points are particularly visible, e.g., in Fig. 5(c).
@@
At small $\gamma$, the slope of the gap increases with increasing $N$, see Figs. 5(a)–(c), which suggests that the weak dissipation ($\gamma\to 0$) and thermodynamic ($N\to\infty$) limits do not commute. Hence, if the thermodynamic limit is taken first, the gap converges to a finite constant of order $J$ (where $J$ is the typical coupling strength of $H$) in the limit of $\gamma\to 0$, signaling a much faster relaxation than expected from perturbation theory. This phenomenon, dubbed anomalous relaxation [120], signals that internal interactions, not external dissipation, govern the relaxation. It has been recently observed in a variety of systems [119, 120, 121, 142, 143, 144, 35, 122, 123, 145, 112], and it is proposed to be a generic feature of ergodic many-body quantum dynamics [121, 122, 123, 146, 147] (via a relation to quantum Ruelle-Pollicott resonances [121, 123]). [^11] The noncommutativity of limits can be understood heuristically as follows [121]. For an $N$-site system with bulk dissipation $\gamma$, there is a natural effective dissipation strength $\gamma_{\mathrm{eff}}=N\gamma$. It is thus manifest that the two limits do not commute: $\lim_{N\to\infty}\lim_{\gamma\to 0}\gamma_{\mathrm{eff}}=0$, while $\lim_{\gamma\to 0}\lim_{N\to\infty}\gamma_{\mathrm{eff}}=\infty$. Nevertheless, the weak-dissipation dynamics of our gamma-matrix model are not fully mixing because of the emergent integrability arising from the fractionalizing into an exponentially large number of gauge sectors with free-fermion excitations. In every sector gauge-equivalent to $\mathfrak{U}=\emptyset$, the Linbdladian has an additional conserved quantity in the limit of $\gamma\to 0$, namely the open gauge strings $W^{\mathrm{ch}}_{\mathcal{P}}$ (together with $H$ itself, these exhaust the conservation laws of the Hamiltonian and hence the $\gamma\to 0$ limit of the Lindbladian). By virtue of these extra conservation laws, there is an infinite number (in the thermodynamic limit) of gaps that close, and the system as a whole evades anomalous relaxation; see, e.g., Fig. 5(d). By choosing the gauge where the particle-number conservation is manifest, we can analytically bound the gap in different particle-number sectors and prove the absence of anomalous relaxation; see Sec. VI.2. Notwithstanding, the (also extensive) number of sectors with finite gaps further illustrates the ubiquity of anomalous relaxation as a generic phenomenon in strongly interacting many-body dynamics.
@@
On the other hand, certain symmetry sectors display the quantum Zeno effect—the closing of the spectral gap with increasing dissipation strength—which was reported for homogeneous dissipation on the square lattice in Ref. 35. For the gap to close in this limit, there must exist an operator $\Sigma_{R}$ that commutes with all jump operators $L_{j}\sim\Gamma_{j}^{\mathrm{ch}}$. Such operators exist in sectors that either are gauge-equivalent to $\mathfrak{U}=\emptyset$ or have $\mathfrak{V}=\emptyset$ and $(-1)^{|\mathfrak{U}|}=-1$. The former will be discussed in detail in Sec. VI.3, and here we focus on the latter. For any choice of $\mathfrak{U}$ with even $|\mathfrak{U}|$, we consider a product $\Sigma_{R}$ of open gauge strings $W_{\mathcal{P}}^{\mu\nu}$ ($\mu,\nu\in\{1,\dots,2k\}$) whose endpoints connect a pair of sites from which a gauge flip emanates, such that all flips are connected by one such string. From Sec. IV.5, $\sigma=\rho_{\bm{w}}\Sigma_{R}$ belongs to the sector labeled by $\mathfrak{U}$ and, because $W_{\mathcal{P}}^{\mu\nu}$ is built from two gamma matrices on each site, it commutes with all chiral elements $\Gamma_{j}^{\mathrm{ch}}$. Hence, $\sigma$ is annihilated by $\mathcal{L}$ in the limit $\gamma_{j}\to\infty$. On the other hand, if $|\mathfrak{U}|$ is odd, one of the gauge flips is not paired up. Hence, there exists a site $j$ for which $\Gamma_{j}^{\mu}$ appears only once in $\Sigma_{R}$, which thus anticommutes with $\Gamma_{j}^{\mathrm{ch}}$, and the gap diverges. These predictions are confirmed in Fig. 5.121212At first sight, this reasoning should imply that the gap diverges as $\gamma\to\infty$ in Fig. 5(d) because $|\mathfrak{U}|$ is odd. However, as discussed in Sec. IV.5, the action of $\Gamma_{j}^{\mathrm{ch}}$ is not to flip all the $\tilde{u}_{ij}$ around a site but to flip the corresponding $v_{j}$. A representative string in the gauge-equivalent representation with $\mathfrak{V}=\emptyset$ (which we are employing in the numerical calculations) is $\Sigma_{R}=D_{j}\Gamma_{j}^{\mathrm{ch}}$. This operator, besides flipping the intralayer gauge fields around site $j$, also anticommutes with the fermionic parity $(-1)^{\mathcal{N}_{f}}$, leading to the result of Fig. 5(d), which is also in agreement with following section, cf. Fig. 6.